OpenStudy (anonymous):

How do I convert y=100*2^x into Log form?

OpenStudy (anonymous):

More importantly, where does the 100 go?

OpenStudy (anonymous):

Sorry, 100*(2^x)

OpenStudy (anonymous):

\[100 \times 2 ^{x}\] =\[\ln \frac{ y }{ 100 } = x \ln 2 . \] \[x = \ln \frac{ y }{ 100 } \div \ln 2 .\]

OpenStudy (anonymous):

divide y by 100

OpenStudy (anonymous):

The end is in log form?

OpenStudy (anonymous):

change the symbol ln to log it won't matter the the base differs from e to 10

OpenStudy (anonymous):

\[\ln = \log_{e} \]

OpenStudy (anonymous):

i took log to both sides or ln

OpenStudy (anonymous):

gotcha, so what was the end in Log, but more importantly, how did you get that?

OpenStudy (anonymous):

So would it be x=Log(2)Y/100?

OpenStudy (anonymous):

\[2^{x}= \frac{ y }{ 100 }\] take the log for both sides \[\log 2 ^{x} = \log \frac{ y }{ 100 }\] and use property that \[\log 2 ^{x} = x \log 2.\] and divide by log2

OpenStudy (anonymous):

man when we write log we mean it is for the base 10

OpenStudy (anonymous):

so the base is constant for all = 10

OpenStudy (anonymous):

you can solve it in another way like \[2^{x} = \frac{ y }{ 100 }\] by definition \[\log_{2} \frac{ y }{ 100} =x \]

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

I understand what I did up until the y/100, can you explain that at all?

OpenStudy (anonymous):

OK let's explain the second method first for the definition of the log \[a ^{b} = c\] thus \[\log_{a}c = b\] and change this symbols into yours.

OpenStudy (anonymous):

where will the 100 go?

OpenStudy (anonymous):

It will stay as it was. why do you want to change it??

OpenStudy (anonymous):

No, I just dont know where it would go because it doesnt look like it fits anywhere? is it part of a? I am confused

OpenStudy (anonymous):

\[a ^{b} = 2^{x} =c = \frac{ y}{ 100 }\] as 100 doesn't have the power x

OpenStudy (anonymous):

So it y just gets divided by it?

OpenStudy (anonymous):

Lets forget about that one. What happens in your original answer after we log both sides?

OpenStudy (anonymous):

i use the property of \[\log a ^{b}= b \log a \]

OpenStudy (anonymous):

So we have log(2)x=xLog(2)

OpenStudy (anonymous):

But is that with a base of 10?

OpenStudy (anonymous):

yea when say just log without bases we mean that the base is 10.My textbook does that always and i have never seen \[\log_{10} \] anyway it won't matter(if you wrote it or not

OpenStudy (anonymous):

So the end is..? I think im getting it

OpenStudy (anonymous):

you can take log with any base put for 10 or e are so common and easier to deal with.You will see ln for both sides in physics a lot as its graph is well-known. write the problem for the start and try to apply what you understand and give full steps.

OpenStudy (anonymous):

So we started with 100*2^x=y and now we have xLog(2)=y/100?

OpenStudy (anonymous):

you took log for both sides not only one ,so log (y/100)

OpenStudy (anonymous):

And that is the final answer with a base of 10?

OpenStudy (anonymous):

@Catch.me

OpenStudy (anonymous):

yea (you can make it any base by constant in both sides)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

you are welcome :D

OpenStudy (anonymous):

x log 2 = Log (y/100) Final answer right? @Catch.me

OpenStudy (anonymous):

right and it would be better if you make x in an isolated term to make it at the final = the answer which is gotten from the definition. \[x= \log_{2} \frac{ y }{ 100 }\] by using the property \[\log_{a} b = \frac{ \log_{c} b }{ \log_{c}a }.\]

OpenStudy (anonymous):

But the base is no longer 10 is it?

OpenStudy (anonymous):

for which equation (write your thinking) c be what ever it is just it has to be constant for both. in your method by taking log for both sides c = 10.

OpenStudy (anonymous):

I just need it to be a base of ten, so im guessing it is all good, thanks for walking me though it!

OpenStudy (anonymous):

so you understood it. so check the properties of the logs there http://tutorial.math.lamar.edu/Extras/AlgebraTrigReview/LogProperties.aspx try yourself first then check it :D